Kim owns storage building that is used exclusively in her business. The building has an adjusted basis of $9,100 (FMV $5,300)I need help with this.
Answer #1

a. Recognized loss on sale = FMV - Basis = 7500 - 9100 - 2200 = (3800)

b. Basis of storage building = 5300 + 2200 = 7500